The apparent magnitude of the sun is - 26.8. Determine its absolute magnitude. It is given that the distance between the sun and the earth is 1.5 x 10^11m.
1
Expert's answer
2019-06-03T10:17:21-0400

The absolute magnitude "M" of a star is the magnitude which the star would have if it were located "d_0=10\text{ pc}" from the observer. If we know the apartment magnitude "m", we can use the following equation to determine the absolute:


"M=m-5\text{lg}\frac{d}{d_0}."

Convert the distance from the Earth to the sun from kilometers to parsecs:


"1.5\cdot10^{-11}\text{ km}\rightarrow4.848\cdot10^{-6}\text{ pc}."

Then


"M=-26.8-5\text{lg}\frac{4.85\cdot10^{-6}}{10}=+4.77."